# Manifolds from Partitions

### 15 Jan 2024 v1

@@ 4. Gallery
Figure 8. The host $G$ is a 4-sphere defined as the join of $C_{15}$ and the Barycentric refinement ${\rm Oct}_{1}$ of the Octahedron $K_{\{2,2,2\}}={\rm Oct}$. The 4-manifold $G$ has $26+15=41$ vertices because ${\rm Oct}_{1}$ has $6+8+12=26$ vertices. The partition complex $P$ is defined by $p=(1,2,3)$. There are already $\sum_{j=0}^{6}{41\choose 6-j}j^{41}$ $=7834446798231431881744478798427378344467982314318817444787984273$ surjective functions from $\{1,\dots,24\}$ onto $\{1,\dots,6\}$. The number of functions from $G$ onto $P$ (as defined in the first page) is even larger as a function must only cover only at least one of the 6 triangles in $P$. Each of these $f$’s produces a 2-manifolds $H(G,f,P)$, colored depending which of the triangles in $P$ is reached. $H(G,f,P)$ is so partitioned into submanifolds with boundary.

### 02 Feb 2024 v2

@@ 5. Gallery
Figure 8. The host $G$ is a 4-sphere defined as the join of $C_{15}$ and the Barycentric refinement ${\rm Oct}_{1}$ of the Octahedron $K_{\{2,2,2\}}={\rm Oct}$. The 4-manifold $G$ has $26+15=41$ vertices because ${\rm Oct}_{1}$ has $6+8+12=26$ vertices. The partition complex $P$ is defined by $p=(1,2,3)$. There are already $\sum_{j=0}^{6}{41\choose 6-j}j^{41}$ $=7834446798231431881744478798427378344467982314318817444787984273$ surjective functions from $\{1,\dots,24\}$ onto $\{1,\dots,6\}$. The number of functions from $G$ onto $P$ (as defined in the first page) is even larger, as a function must only cover only at least one of the 6 triangles in $P$. Each of these $f$’s produces a 2-manifolds $H(G,f,P)$, colored depending which of the triangles in $P$ is reached. $H(G,f,P)$ is so partitioned into sub-manifolds with boundary.